1.) Describe how you prepare 100 mL of an aqueous 0.05 M KOH solution from a stock solution with a concentration of of 25.0 g/L. You have 2.0 L of this stock solution. Show the calculation and write up the preparation of the dilute solution in the format of the Experimental section of a publication.

2.) You have a 1.0 L standard solution of a 500 ppm copper solution. Prepare a 20 mL solution of 10 ppm.
Answer #1

1) First we need to calculate molar concentration of stock solution.

We know that, Molarity = No. of moles of solute / Volume of solution in L

Here, Volume of solution = 1 L

We have, No. of moles of KOH = Mass / Molar Mass

Molar MAss of KOH = 39.09 + 16.00 + 1.0079 = 56.10 g / mol

Therefore, No. of moles of KOH = 25.0 g / 56.10 g/mol = 0.446 mol

Hence, Molarity of KOH solution = 0.446 mol / 1 L = 0.446 M

Now, we can use dilution formula to find out volume of stock solution required to prepare 0.05 M KOH solution.

We have, M stock x V stock = M dilute x V dilute

Therefore, V stock = M dilute x V dilute / M stock = 0.05 M x 100 ml / 0.446 M = 11.2 ml

Procedure : Take 11.2 ml of stock solution into 100 ml volumetric flask and dilute up to the mark with distilled water.

2) Volume of stock solution required to prepare 20 ml solution of 10 ppm copper is

V stock = M dilute x V dilute / M stock =10 ppm x 20 ml / 500 ppm = 0.4 ml

Take 0.4 ml of 500 ppm solution of Copper and dilute it to 20 ml using distilled water.